When unpolarised light is incident on a glass plate at a particular angle, it is observed that the reflected beam is linearly polarised. what is the angle of the refracted beam with respect to the normal? Refractive index of glass is 1.52

Please explain in detail and write all relevant formulas. Thank you.
Answer #1

For a reflected ray to be completely polarised the angle of incidence is called angle of polarisation.Now according to Brewster's law,

If
i = the angle of incident ,

p = the angle of polarisation

r = the angle of refraction

Rg = the refractive index of the glass

we have ,

   tan p = 1.52

   but here
   p = i
tan i = 1.52

so,
i = tan-1 ( 1.52 )

i = 56.65o

Thus, angle of incident is :

   i = 56.65o

we have another relation as :

angle of refraction + angle of polarization = 90 deg.

i.e    r + i = 90
Hence ,

angle of refraction = 90-angle of polarization

   = 90 - 56.650


therefore, angle of refraction:

  r = 90 - 58.31 = 33.34 o

angle of refraction of the beam = r = 31.69 o